ed is 7 times as old as laura. if he is 48 years older than she is, how old is she?
Answer by stanbon(75887)

ed is 7 times as old as laura. if he is 48 years older than she is, how old is she?
--------
E = 7*L
E = L + 48
-------------------
7L = L + 48
6L = 48
Laura = 8 yrs
-----------------
Cheers,
Stan H.
------------